The probability that a football player weighs more than 230 pounds is .69, that he is at
least 75 inches tall is .55, and that he weights more than 230 pounds and is at least 75 inches tall is .43. Find the probability that a randomly selected football player weighs more than 230 pounds or is at least 75 inches tall.

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

p(>230)=0.69
P(>=75)=0.43
P(BOTH)=0.69*0.43 if independent
P(OR) is P(A)+p(B)-2P(both A and B), the last term because of double counting
=0.69+0.43-0.2967, and round to .30
=1.12-0.30=0.82

Answer by ikleyn(52781)

.

            The correct solution is as follows.


Use the general formula  of the Probability theory


          P(A U B) = P(A) + P(B) - P(A ∩ B).


In your case P(A) = 0.69  (weighs more than 230 pounds);

             P(B) = 0.55  (is at least 75 inches tall),

             P(A ∩ B) = 0.43  (weights more than 230 pounds and is at least 75 inches tall).



Substitute the given data into the formula and calculate


         P(A U B) = 0.69 + 0.55 - 0.43 = 0.81.      ANSWER

Solved (correctly).

--------------

The solution by @Boreal is INCORRECT.